zackariahharriett:

The steps below show the incomplete solution to find the value of x for the equation 6x − 2x − 5 = −6 + 21: Step 1: 6x − 2x − 5 = −6 + 21 Step 2: 6x − 2x − 5 = 15 Step 3: 4x − 5 = 15 Which of these is most likely the next step? (4 points) 4x = 10 4x = 3 4x = 45 4x = 20

sophia333:

4x-5=-6+21 4x - 5= 15 4x= 20 ÷4 x= 5

sophia333:

This make sense?

zackariahharriett:

thank u

sophia333:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @zackariahharriett thank u \(\color{#0cbb34}{\text{End of Quote}}\) No problem :)